# The Energy Cost of Artificial Intelligence Lifecycle in Communication Networks

### v3

@@ VI Energy Cost of Training
To train a model, training data and a machine learning techniques is needed. In this study, we consider neural network architectures, centralized offline training, however, extensions of this work could consider online, distributed, federated or transfer approaches as well. As discussed in Section V, the training data typically goes through a preprocessing phase and then split into training and evaluation sets, with a $\beta$ split ratio as:
@@
where $N_{\mathrm{S,T}}$ and $N_{\mathrm{S,E}}$ represent the number of samples for training and evaluation, respectively. While many neural network architectures are available for AI models development, in this work, we focus on four representative types. Namely, Multilayer Perceptrons [11], Convolutional Neural Networks [38], Kolmogorov–Arnold Networks [39], and transformers [40]. Training a neural network involves two fundamental phases: forward propagation and backward propagation, depicted with blue and red in Fig. 6, respectively, on an example MLP. During the forward propagation, input data passes through the network layers. Each layer performs specific tasks, such as computing a linear combination of weights and biases, followed by applying activation functions in dense (fully connected) layers. Convolutional layers perform convolutional operations, while pooling mechanisms are applied on pooling layers. At the end of this phase, the loss is calculated by comparing the predicted outputs to the actual target values. Subsequently, the model performs backward propagation, where it calculates gradients of the loss with respect to each parameter, updates these gradients in reverse through the layers, and adjusts the weights and biases accordingly. This process completes one epoch. Consequently, we can calculate the energy costs associated with forward and backward propagation, leading to an understanding of the overall energy consumption for training, model evaluation, and subsequent inference. In the following sections, we showcase the computational complexity of training and evaluating the aforementioned Neural Networks.
@@ VI-A Computational Complexity of MLP, CNN, KAN and transformer
where the first term in the sum represents one multiplication and one addition ($=2$ FLOPs) related to the weight and bias corresponding to the edges times the number of nodes in that layer ($=M_{l-1}M_{l}$), while the second term represents the summation of all the values of the incoming edges and the application of the activation function ($=2$ FLOPs) times the number of nodes in that layer ($M_{l}$).
@@
where $K_{\mathrm{r}}$ and $K_{\mathrm{c}}$ are the kernel (filter) height and width. Moreover, $P_{\mathrm{r}}$ and $P_{\mathrm{c}}$ are the padding sizes for height and width. Furthermore, $S_{\mathrm{r}}$ and $S_{\mathrm{c}}$ are the stride values for height and width. Finally, $N_{\mathrm{f}}$ is the number of filters in the layer. For down-sampling the input tensor data, a pooling layer is needed, and the number of FLOPs of such layer is expressed as:
@@
KANs
$M$ =${}_{\mathrm{KAN}}=\sum_{l=1}^{L-1}(\underbrace{M_{\mathrm{NLF}}M_{l-1}}_{\text{From B-spline}}$ (16)
where $M_{\mathrm{NLF}}$ denote as the number of FLOPs contributed from the B-spline activation function across all input elements. Additionally, the second term accounts for the computational costs arising from the combination of input and output dimensions ($M_{l-1}$ and $M_{l}$) with the operations performed by the B-spline transformation. These operations include the iterative evaluation of spline basis functions and the associated transformations, governed by the spline order ($K$) and the number of intervals in the grid ($G$). The total FLOPs of a KANs network is calculated as the summary of all KAN layers similarly as in the case of MLPs in Eq. 12, where total $L$ layers can be expressed as:
$M_{\mathrm{KAN,FP}}=$ $N_{\mathrm{epochs}}N_{\mathrm{S,T}}M_{\text{KAN}}.$ (17)
@@
Based on the attention mechanism, the transformers [40] represent a paradigm shift in deep learning, replacing traditional architectures such as CNN with self-attention mechanisms. While enabling significant breakthroughs in AI through efficient parallel processing and robust modeling of long-range dependencies, they are relatively complex. In addition to attention layers, the transformer architecture also includes MLP, normalization, and positional encoding, making the analytic expression for the FLOP computation intractable. Following existing literature [43] and approximations  [^4] https://www.gaohongnan.com/playbook/training/how_to_calculate_flops_in_transformer_based_models.html, we approach the estimation as follows: $M_{\text{ATT}}=2(\underbrace{CN_{\text{embed}}3N_{\text{embed}}}_{\text{K, Q, V positional embedding}}+\underbrace{C^{2}N_{\text{embed}}}_{\text{Attention scores}}$ (18) $+\underbrace{N_{\text{head}}C^{2}N_{\text{embed}}/N_{\text{head}}}_{\text{reduce}}+\underbrace{CN_{\text{embed}}N_{\text{embed}}}_{\text{Projection}}),$ where $C$ is the context length, $N_{\text{embed}}$ is the size of the embedding and $N_{\text{head}}$ is the number of heads. The complexity of the decoder in a transformer is then given by:
@@
where $N_{\text{decoder\_blocks}}$ and $FFS$ stand for the number of decoder blocks and the feed-forward size.
@@
VI-B Model Training
To determine the computational complexity of training a model, backward propagation needs to also be considered. The standard approximation, also confirmed in [44], is that the backward propagation requires two times the floating-point operations per cycle per core (FLOPs) of the forward propagation. In particular, backward propagation requires matrix multiplications for updating the weights and propagating the gradient. Thus, we can approximate the computing complexity of training a model as:
@@ VI-B Model Training
We can obtain the energy consumption of the entire training process as:
@@
From Eq. 12151720, and 22 we can observe that the training data set size scales linearly with energy consumption of the model training, as it does not affect the cost of single forward/backward pass. On the other hand, training the same model on a state-of-the-art GPU with a processing power of $312\leavevmode\nobreak\ [TFLOPs/s]$ and power consumption of $400\leavevmode\nobreak\ [W]$ is approximately $28$ times more energy efficient than using a CPU, which operates at $13.824\leavevmode\nobreak\ [TFLOPs/s]$ and consumes $500\leavevmode\nobreak\ [W]$.
@@
Fig. 7 illustrates the relationship between the energy consumption of a model and two architectural parameters: layers and nodes in a layer. Naturally, as the architectures of the models become deeper and wider, the computational complexity also increases, and this leads to higher energy consumption. For instance, in the case where $M_{l}=10$ and $L=5$, the energy consumption of the transformer, KAN, and CNN models are approximately $173$, $149$, and $81$ times greater than that of the MLP model. However, some architectures, such as KANs, are known to perform better with fewer layers  [39], therefore it is unlikely to find a KAN with $L>5$. While at a first glance, it seems that MLPs may be significantly more energy efficient than KANs, the actual difference for the same performance will be less prominent for many applications. CNNs, and especially transformers are known to be computationally demanding and subsequently exhibit high energy consumption, while at the same time larger architectures proved better performance in various application areas.
@@ VI-C Model Evaluation
The model evaluation starts once training is complete. This functionality tests the performance of the model on a separate dataset. During the evaluation, the model processes the data using forward propagation without any adjustments to its parameters. Therefore, the energy consumption and energy consumption of model evaluation is expressed as:
@@
It is worth noting that the total energy consumption during training and evaluation is significantly influenced by the chosen evaluation strategy. While a simple train-test split requires training the model only once, more robust techniques like $k$-fold cross-validation necessitate $k$ complete training cycles. For $k$-fold cross-validation, the total training energy consumption becomes $k$ times the value given in Eq. 22, as the model must be retrained from scratch for each fold. Similarly, techniques such as nested cross-validation or repeated $k$-fold cross-validation further multiply the required training computations and corresponding energy costs, in which we have also incorporated $k$-fold cross-validation considerations into our developed tool.
@@ VII Energy Cost of Inference
The complexity of making an inference depends on the number of FLOPs required for forward propagating with input sample size $N_{\mathrm{I,P}}$, i.e., $N_{\mathrm{inf}}=M_{\mathrm{model}}N_{\mathrm{I,P}}$. For example, when we adopt the aforementioned MLP model with $N_{\mathrm{I,P}}=51$, the complexity is calculated as $N_{\mathrm{inf}}=23651=12,036$ FLOPs. Therefore, the energy consumption of the forward propagation and the corresponding inference can be calculated as:
@@
We can then observe from Eq. 23 and 24 that the cost of inference is the same as in evaluation when $N_{\mathrm{S,E}}$ is equal to $N_{\mathrm{I,P}}$. Thus, we can conclude that the computational complexity between finishing one training and inference is on the magnitude of $3N_{\mathrm{epochs}}N_{\mathrm{S,T}}/N_{\mathrm{I,P}}$. This indicates that the training process involves significantly more computational operations, particularly when the number of epochs or $N_{\mathrm{S,T}}>>N_{\mathrm{I,P}}$. On the other hand, the energy consumption per bit depends on the model complexity and hardware capability. Moreover, the factor of three in the energy consumption per bit during training versus inference highlights the additional computational load inherent to training.
@@
Table III: Total energy consumption of developing ($E_{\mathrm{D}}$) for different models, with contributions of individual data-manipulating components.
Component MLP CNN KAN Transformer
$E_{\mathrm{DC}}\leavevmode\nobreak\ [J]$ $8.44\times 10^{-2}$ ($89.12\%$) $8.44\times 10^{-2}$ ($12.37\%$) $8.44\times 10^{-2}$ ($6.31\%$) $8.44\times 10^{-2}$ ($3.72\%$)
$E_{\mathrm{pre}}\leavevmode\nobreak\ [J]$ $1.54\times 10^{-6}$ ($0.00\%$) $1.54\times 10^{-6}$ ($0.00\%$) $1.54\times 10^{-6}$ ($0.00\%$) $1.54\times 10^{-6}$ ($0.00\%$)
$E_{\mathrm{train}}\leavevmode\nobreak\ [J]$ $1.03\times 10^{-2}$ ($10.86\%$) $5.97\times 10^{-1}$ ($87.48\%$) $1.25$ ($93.53\%$) $2.18$ ($96.12\%$)
$E_{\mathrm{eval}}\leavevmode\nobreak\ [J]$ $1.72\times 10^{-5}$ ($0.02\%$) $9.95\times 10^{-4}$ ($0.15\%$) $2.08\times 10^{-3}$ ($0.16\%$) $3.64\times 10^{-3}$ ($0.16\%$)
$E_{\mathrm{D}}\leavevmode\nobreak\ [J]$ $9.47\times 10^{-2}$ $6.83\times 10^{-1}$ $1.34$ $2.27$
In this section, we describe the end-to-end energy consumption of the AI model lifecycle for making inferences with the aforementioned neural network architectures. First, we consider the energy consumption involved in developing the model, as illustrated on the left in Fig. 1, which can be expressed as follows:
$E_{\mathrm{D}}=E_{\mathrm{DC}}+E_{\mathrm{pre}}+E_{\mathrm{train}}+E_{\mathrm{eval}},$ (25)
and the energy consumption per bit that can be expressed as:
@@
To compare the energy consumption of all the data manipulation components required for developing and deploying an AI-enhanced system, we assume that the models are trained and evaluated using 5-fold cross-validation (as discussed in Section VI). The models require $256$ samples with an input size of $10$, i.e., $N_{\mathrm{S}}I_{\mathrm{S}}=2560$, collected from a device over a wireless network configured with the following parameters: $P_{\mathrm{T}}=200\leavevmode\nobreak\ [mW]$, $P_{\mathrm{R}}=5\leavevmode\nobreak\ [mW]$, $R_{\mathrm{T}}=R_{\mathrm{R}}=10\leavevmode\nobreak\ [Mbps]$, $N_{\mathrm{\_,cycle},l}=100\leavevmode\nobreak\ [c/b]$, and $P_{\mathrm{\_,cycle}}=10^{-10}\leavevmode\nobreak\ [W/c]$. We also assume the link between the device and the server is stable, and there is no retransmission needed. Each layer introduces DP and CP overhead of $10\%$ and $5\%$, respectively. Furthermore, normalization is utilized as the data transformation procedure, as discussed in Section V. Finally, model training requires $10$ epochs ($N_{\mathrm{epochs}}=10$) and utilizes $5$-fold validation. The energy consumption of this example is shown in Table III and reveals the contribution of each component to the total energy cost of developing models. Specifically, data collection dominates the total energy cost for developing a very simple model based on a 3-layer, 10-nodes-per-layer MLP architecture. For the same parameters, the relative contribution of the data collection decreases for the KAN, CNN and transformer architectures, making the energy cost of training the primary contributor. The eCAL calculator enables changing the hyperparameters of the four architectures while also adding well known CNN architectures such as ResNets [45], VGGs [46] or transformer-based such as Baichuan 2 [47].
@@
Once the model is developed, it is packaged and deployed in the communication system, where it can produce inference in the form of continuous or discrete outputs. During its operational lifecycle, the model will be presented with some input data and requested to produce the corresponding inference. Thus, we can express the associated energy cost as follows:
@@
Next, we calculate the total energy cost over the lifetime of an AI model in the system, also referred to as $eCAL_{\mathrm{abs}}$ (measured in [$J$]). It is defined as the sum of energy consumed during the model development, $E_{\mathrm{D}}$ and the energy required for performing an inference $E_{\mathrm{inf,p}}$, multiplied by the number of times $\gamma$ the inference is performed with the currently deployed model. In addition, as edge cloud systems may use virtualization in developing and operating the models, we enable considering the impact of such technologies through a factor. As a result, $eCAL_{\mathrm{abs}}$ can be expressed as:
@@
where $\gamma_{\mathrm{v}}$ represents the contribution from the virtualization, and $\gamma\in\left[0,\infty\right)$.
@@
Figure 9: Energy cost of AI model lifecycle ($eCAL$) for different models over the number of inferences ($\gamma$), log scale on both axes.
Finally, the proposed metric computing the corresponding energy consumption per bit over the lifecycle of the currently deployed AI model can be expressed as:

### v4

@@ VI Energy Cost of Training
To train a model, training data and machine learning techniques are needed. In this study, we consider neural network architectures with centralized offline training, however, extensions of this work could consider online, distributed, federated or transfer approaches as well. As discussed in Section V, the training data typically goes through a preprocessing phase and is then split into training and evaluation sets, with a $\beta$ split ratio as:
@@
where $N_{\mathrm{S,T}}$ and $N_{\mathrm{S,E}}$ represent the number of samples for training and evaluation, respectively. For studying the energy cost of training, in this paper we focus on four representative neural network architectures [^4] Note that the implementation of the calculator supports many other architectures using pytorch., namely, Multilayer Perceptrons [11], Convolutional Neural Networks [44], Kolmogorov–Arnold Networks [45], and transformers [46]. MLPs established the foundations of deep learning as universal function approximators, enabling the first wave of practical neural network models. CNNs marked a breakthrough by exploiting spatial hierarchies, revolutionizing computer vision and inspiring domain-specific architectures. Transformers represent the most recent paradigm shift, introducing attention mechanisms that now dominate language processing through the GPT models. Finally, KANs are included as a novel, explainable architecture compared to MLPs.
@@
Training a neural network involves two fundamental phases: forward propagation and backward propagation, depicted with blue and red arrows in Fig. 7, respectively, on an example MLP. During the forward propagation, input data passes through the network layers. Each layer performs specific tasks, such as computing a linear combination of weights and biases, followed by applying activation functions in dense (fully connected) layers. Convolutional layers perform convolutional operations, while pooling mechanisms are applied on pooling layers. At the end of this phase, the loss is calculated by comparing the predicted outputs to the actual target values. Subsequently, the model performs backward propagation, where it calculates gradients of the loss with respect to each parameter, updates these gradients in reverse order through the layers, and adjusts the weights and biases accordingly. This process completes one epoch. Consequently, we can calculate the energy costs associated with forward and backward propagation, leading to an understanding of the overall energy consumption for training, model evaluation, and subsequent inference. In the following subsections, we showcase the computational complexity of training and evaluating the aforementioned Neural Networks.
@@ VI-A Computational Complexity of MLP, CNN, KAN and transformer
where the first term in the sum represents one multiplication and one addition (=2=2 FLOPs) related to the weight and bias corresponding to the edges times the number of nodes in that layer ($M_{l-1}M_{l}$), while the second term represents the summation of all the values of the incoming edges and the application of the activation function (=2=2 FLOPs) times the number of nodes in that layer ($M_{l}$).
@@
where $K_{\mathrm{r}}$ and $K_{\mathrm{c}}$ are the kernel (filter) height and width, $P_{\mathrm{r}}$ and $P_{\mathrm{c}}$ are the padding sizes for height and width, $S_{\mathrm{r}}$ and $S_{\mathrm{c}}$ are the stride values for height and width, and $N_{\mathrm{f}}$ is the number of filters in the layer. For down-sampling the input tensor data, a pooling layer is needed, and the number of FLOPs of such layer is expressed as:
@@
KANs
$M$ ${}_{\mathrm{KAN}}=\sum_{l=1}^{L-1}(\underbrace{M_{\mathrm{NLF}}M_{l-1}}_{\text{From B-spline}}$ (17)
where $M_{\mathrm{NLF}}$ denote as the number of FLOPs contributed from the B-spline activation function across all input elements. The second term accounts for the computational costs arising from the combination of input and output dimensions ($M_{l-1}$ and $M_{l}$) with the operations performed by the B-spline transformation. These operations include the iterative evaluation of spline basis functions and the associated transformations, governed by the spline order ($K$) and the number of intervals in the grid ($G$). The total FLOPs of a KANs network is calculated as the sum of all KAN layers similarly as in the case of MLPs in Eq. (13), where total $L$ layers can be expressed as:
$M_{\mathrm{KAN,FP}}=$ $N_{\mathrm{epochs}}N_{\mathrm{S,T}}M_{\text{KAN}}.$ (18)
@@
Based on the attention mechanism, the transformers [46] represent a paradigm shift in deep learning, replacing traditional architectures such as CNN with self-attention mechanisms. While enabling significant breakthroughs in AI through efficient parallel processing and robust modeling of long-range dependencies, they are relatively complex. In addition to attention layers, the transformer architecture also includes MLP, normalization, and positional encoding, making the analytic expression for the FLOP computation intractable. Following existing literature [49] and approximations [^5] https://www.gaohongnan.com/playbook/training/how_to_calculate_flops_in_transformer_based_models.html, we approach the estimation as follows:
$M_{\text{ATT}}=2(\underbrace{CN_{\text{embed}}3N_{\text{embed}}}_{\text{K, Q, V positional embedding}}+\underbrace{C^{2}N_{\text{embed}}}_{\text{Attention scores}}$ (19)
$+\underbrace{N_{\text{head}}C^{2}N_{\text{embed}}/N_{\text{head}}}_{\text{reduce}}+\underbrace{CN_{\text{embed}}N_{\text{embed}}}_{\text{Projection}}),$
where $C$ is the context length, $N_{\text{embed}}$ is the size of the embedding and $N_{\text{head}}$ is the number of heads. The complexity of the decoder in a transformer is then given by:
@@
where $N_{\text{decoder\_blocks}}$ and $FFS$ stand for the number of decoder blocks and the feed-forward size, respectively. The total complexity of forward propagation is:
@@
VI-B Energy Estimation for the Model Training
To determine the computational complexity of training a model, backward propagation needs to also be considered. The standard approximation, also confirmed in [50], is that the backward propagation requires two times the floating-point operations per cycle per core (FLOPs) of the forward propagation. In particular, backward propagation requires matrix multiplications for updating the weights and propagating the gradient. Thus, we can approximate the computing complexity of training a model with Eq. 22 and obtain the energy consumption of the entire training process with E1. 23.
@@ VI-B Energy Estimation for the Model Training
From Eqs. (13), (16), (18), (21), and (23), we can observe that the training data set size scales linearly with energy consumption of the model training, as it does not affect the cost of a single forward/backward pass. On the other hand, training the same model on a state-of-the-art GPU with a processing power of $312\penalty 10000\ [TFLOPs/s]$ and power consumption of $400\penalty 10000\ [W]$ is approximately $28$ times more energy efficient than using a CPU, which operates at $13.824\penalty 10000\ [TFLOPs/s]$ and consumes $500\penalty 10000\ [W]$.
Numerical and Empirical Validation
@@
Fig. 8 illustrates the relationship between the energy consumption of a model and two architectural parameters: the number of layers and the number of nodes in a layer. As the architectures of the models become deeper and wider, the computational complexity also increases, and this leads to higher energy consumption. For instance, in the case where $M_{l}=10$ and $L=5$, the energy consumption of the transformer, KAN, and CNN models are approximately $173$, $149$, and $81$ times greater than that of the MLP model. However, some architectures, such as KANs, are known to perform better with fewer layers [45], therefore it is unlikely to find a KAN with $L>5$. While at a first glance, it seems that MLPs may be significantly more energy efficient than KANs, the actual difference for the same performance will be less prominent for many applications. CNNs, and especially transformers are known to be computationally demanding and subsequently exhibit high energy consumption, while at the same time larger architectures proved better performance in various application areas.
Figure 9: Training energy consumption benchmarks for the selected fundamental neural architectures.
Fig. 9 benchmarks the model training energy required by eCAL against (a) a simple fundamental physics approach $E=P_{avg}\cdot t$ that considers the average power consumption of the GPUs over the training time, (b) the open-source calflops666calflops,https://pypi.org/project/calflops/0.0.2/ library able to compute the theoretical amount of FLOPs for neural networks implemented in pytorch, (c) Code Carbon [^7] Code Carbon, https://github.com/mlco2/codecarbon, and (d) a worst case scenario that assumes the GPU is used at maximum capacity during the entire training time computed as $E=P_{max}\cdot t$. For eCAL as per Eq. (23), calflops and worst case, we set the value of $P_{max}=10W$ as typical for an Apple M2 GPU. For the fundamental physics we measured the average power used by the GPU at $P_{avg}=0.38W$ [^8] This study can be replicated on other set-ups with https://github.com/sensorlab/eCAL/blob/main/scripts/ All_Theoretical_Empirical_Training_Energy.py using the PowerMetrics tool. Code Carbon outputs energy, power and time duration measurements. It measures the power and time duration, plotted as Code Carbon ($P\cdot t$), through system tools, such as PowerMetrics for Apple silicon or a dedicated python library for NVIDIA GPUs, while it computes the energy, plotted as Code Carbon E, based on a model. The results show that eCAL and calflops provide very similar energy footprint estimated for MLPs, CNN and Transformers based networks. For MLPs, the difference is negligible, for CNNs eCAL slightly overestimates compared to calflops, while for transformers it slightly underestimates. The slight differences are due to the fact that it is difficult to harmonize all hyperparameters between the theoretical expressions that reflect the neural network architectures and corresponding software implementation that instantiates pytorch models that calfplos relies on. For KANs as newer architectures, calflops underestimates compared to the eCAL due to the fact that the implementation of the architecture relies on several optimizations that reduce the number of FLOPs compared to the theoretical expressions.
The results also show that assuming average power usage during the entire training cycle with the fundamental physics expression $E=P_{avg}\cdot t$ overestimates the consumption. The overestimation is more prominent in simple examples used for this study where the data loading and model saving during training take a significant part of the training time when the GPU is not active. However, when these steps become negligible, this estimation will close in to eCAL and calflops showing the validity of our approach. Code Carbon E measures significantly less consumed energy while Code Carbon ($P\cdot t$) measures significantly more than eCAL and calflops. These results also confirm the existing challenges regarding measuring energy and CO2 footprint of software [51] and confirm the timeliness of our work.
@@ VI-C Model Evaluation
The model evaluation starts once training is complete. This functionality tests the performance of the model on a separate dataset. During the evaluation, the model processes the data using forward propagation without any adjustments to its parameters. Therefore, the energy consumption of model evaluation is expressed as:
@@
It is worth noting that the total energy consumption during training and evaluation is significantly influenced by the chosen evaluation strategy. While a simple train-test split requires training the model only once, more robust techniques like $k$-fold cross-validation necessitate $k$ complete training cycles. For $k$-fold cross-validation, the total training energy consumption becomes $k$ times the value given in Eq. (23), as the model must be retrained from scratch for each fold. Similarly, techniques such as nested cross-validation or repeated $k$-fold cross-validation further multiply the required training computations and corresponding energy costs, in which we have also incorporated $k$-fold cross-validation considerations into eCAL calculator.
Figure 10: Model performance (higher $R^{2}$ is better) vs. computational cost (GFLOPs) for five different architectures across the three datasets. The dashed lines indicate the Pareto front.
Table IV: Total energy consumption of developing ($E_{\mathrm{D}}$) for different models, with contributions of individual data-manipulating components.
Component MLP CNN KAN Transformer
$E_{\mathrm{DC}}\penalty 10000\ [J]$ $8.44\times 10^{-2}$ ($89.12\%$) $8.44\times 10^{-2}$ ($12.37\%$) $8.44\times 10^{-2}$ ($6.31\%$) $8.44\times 10^{-2}$ ($3.72\%$)
$E_{\mathrm{pre}}\penalty 10000\ [J]$ $1.54\times 10^{-6}$ ($0.00\%$) $1.54\times 10^{-6}$ ($0.00\%$) $1.54\times 10^{-6}$ ($0.00\%$) $1.54\times 10^{-6}$ ($0.00\%$)
$E_{\mathrm{train}}\penalty 10000\ [J]$ $1.03\times 10^{-2}$ ($10.86\%$) $5.97\times 10^{-1}$ ($87.48\%$) $1.25$ ($93.53\%$) $2.18$ ($96.12\%$)
$E_{\mathrm{eval}}\penalty 10000\ [J]$ $1.72\times 10^{-5}$ ($0.02\%$) $9.95\times 10^{-4}$ ($0.15\%$) $2.08\times 10^{-3}$ ($0.16\%$) $3.64\times 10^{-3}$ ($0.16\%$)
$E_{\mathrm{D}}\penalty 10000\ [J]$ $9.47\times 10^{-2}$ $6.83\times 10^{-1}$ $1.34$ $2.27$
To illustrate the trade-off between model performance and computational complexity, and to demonstrate how eCAL supports Pareto-front analyses, we evaluate three wireless datasets (Wi-Fi, Bluetooth Low Energy (BLE), and cellular), four fundamental architectures from Section VI, and a ResNet combining CNN and MLP blocks—a configuration known to perform well for wireless localization [52]. Fig. 10 shows model performance ($R^{2}$) versus computational complexity at inference time $M_{\mathrm{model,FP}}$. For the univariate BLE dataset, the highest $R^{2}$ scores are achieved by MLP and KAN models forming the Pareto front, while CNN, ResNet, and Transformer models underperform with higher computational costs. Similar trends appear for the 1D multivariate UMU dataset, where the Transformer performs poorly ($R^{2}=-1$) and is notably outperformed by simpler MLPs. With more signals (multivariate data), CNN performance improves, and these models appear on the Pareto front. On the larger, more complex multivariate CTW dataset, which includes 2D antenna array shapes, Transformer-based models perform better and also lie on the Pareto front. Overall, these results show that more complex neural architectures do not necessarily yield better performance. The balance between model complexity and accuracy depends on both the architecture and the nature and quality of the data.
@@ VII Energy Cost of Inference
The complexity of making an inference depends on the number of FLOPs required for forward propagating with input sample size $N_{\mathrm{I,P}}$, i.e., $N_{\mathrm{inf}}=M_{\mathrm{model}}N_{\mathrm{I,P}}$. For example, when we adopt the aforementioned MLP model with $N_{\mathrm{I,P}}=51$, the complexity is calculated as $N_{\mathrm{inf}}=236\times 51=12,036$ FLOPs. The energy consumption of the forward propagation and the corresponding inference can be calculated as:
@@
We can then observe from Eqs. (24) and (25) that the cost of inference is the same as in evaluation when $N_{\mathrm{S,E}}$ is equal to $N_{\mathrm{I,P}}$. Thus, we can conclude that the computational complexity between finishing one training and inference is on the magnitude of $3N_{\mathrm{epochs}}N_{\mathrm{S,T}}/N_{\mathrm{I,P}}$. This indicates that the training process involves significantly more computational operations, particularly when the number of epochs or $N_{\mathrm{S,T}}>>N_{\mathrm{I,P}}$. On the other hand, the energy consumption per bit depends on the model complexity and hardware capability. Moreover, the factor of three in the energy consumption per bit during training versus inference highlights the additional computational load inherent to training.
@@
In this section, we describe the end-to-end energy consumption of the AI model lifecycle for making inferences with the aforementioned neural network architectures. First, we consider the energy consumption involved in developing the model, as illustrated on the left in Fig. 1, which can be expressed as in Eq. (26) with the energy consumption per bit expressed in Eq. (27).
$E_{\mathrm{D}}=E_{\mathrm{DC}}+E_{\mathrm{pre}}+E_{\mathrm{train}}+E_{\mathrm{eval}}.$ (26)
@@
To compare the energy consumption of all the data manipulation components required for developing and deploying an AI-enhanced system, we assume that the models are trained and evaluated using 5-fold cross-validation (as discussed in Section VI). The models require $256$ samples with an input size of $10$, i.e., $N_{\mathrm{S}}I_{\mathrm{S}}=2560$, collected from a device over a wireless network configured with the following parameters: $P_{\mathrm{T}}=200\penalty 10000\ [mW]$, $P_{\mathrm{R}}=5\penalty 10000\ [mW]$, $R_{\mathrm{T}}=R_{\mathrm{R}}=10\penalty 10000\ [Mbps]$, $N_{\mathrm{\_,cycle},l}=100\penalty 10000\ [c/b]$, and $P_{\mathrm{\_,cycle}}=10^{-10}\penalty 10000\ [W/c]$. We also assume the link between the device and the server is stable, and there is no retransmission needed. Each layer introduces DP and CP overhead of $10\%$ and $5\%$, respectively. Furthermore, normalization is utilized as the data transformation procedure, as discussed in Section V. Finally, model training requires $10$ epochs ($N_{\mathrm{epochs}}=10$) and utilizes $5$-fold validation. The energy consumption of this example is shown in Table IV and reveals the contribution of each component to the total energy cost of developing models. Specifically, data collection dominates the total energy cost for developing a very simple model based on a 3-layer, 10-nodes-per-layer MLP architecture. For the same parameters, the relative contribution of the data collection decreases for the KAN, CNN and transformer architectures, making the energy cost of training the primary contributor. The eCAL calculator enables changing the hyperparameters of the four architectures while also adding well known CNN architectures such as ResNets [53], VGGs [54] or transformer-based such as Baichuan 2 [55].
To complement Table IV and provide a more comprehensive energy profile beyond the analysis in Section IV, Fig. 11 shows the energy consumption breakdown that also includes wired portions of the network including fronthaul and backhaul, by using the microcell splits reported in [56]. We examine three fronthaul/backhaul combinations, i.e. Fiber/Fiber, Copper/Radio, and Copper/Fiber to illustrate how the energy mix shifts when wired paths dominate. For computationally light models such as a relatively shallow MLP, deployment energy is no longer dominated by radio access network (RAN)-based data collection as in Table IV; the wired infrastructure becomes the primary contributor, reaching roughly 90% of the total in the Fiber/Fiber case. For complex models such as the simple transformer considered in this work, training remains the largest single component in energy consumption, but its share drops markedly once wired costs are included, which falling from over 96% in the wireless-only view from Table IV to approximately 68–78% across the three scenarios as depicted in Fig. 11. Importantly, the underlying absolute energies (in Joules) are unchanged: we add fronthaul and backhaul (derived from the microcell splits) to the per-model totals and then recompute percentage shares. The RAN data-collection energy is exactly the same as in Table IV. Taken together, these results show a holistic energy assessment in modern hybrid networks, and emphasize the flexibility of the eCAL framework in modeling diverse scenarios.
Figure 11: Comparison of energy consumption of training ($E_{\mathrm{train}}$, highlighted in upper part of the bars) and data collection ($E_{\mathrm{DC}}$, highlighted in lower part of the bars, divided into three parts: RAN, Fronthaul, and Backhaul for different models in different Fronthaul/Backhaul combinations.
@@
Once the model is developed, it is packaged and deployed in the communication system, where it can produce inference in the form of continuous or discrete outputs. During its operational lifecycle, the model will be presented with input data and requested to produce the corresponding inference. Thus, we can express the associated energy cost as follows:
@@
Next, we calculate the total energy cost over the lifetime of an AI model in the system, also referred to as $eCAL_{\mathrm{abs}}$ (measured in [$J$]). It is defined as the sum of energy consumed during the model development, $E_{\mathrm{D}}$ and the energy required for performing an inference $E_{\mathrm{inf,p}}$, multiplied by the number of times $\gamma\in\left[0,\infty\right)$ the inference is performed with the currently deployed model. In addition, as edge cloud systems may use virtualization in developing and operating the models, we enable considering the impact of such technologies through an additional factor. As a result, $eCAL_{\mathrm{abs}}$ can be expressed as:
@@
where $\gamma_{\mathrm{v}}$ represents the contribution from the virtualization, with $\gamma_{\mathrm{v}}=0$ corresponding to no virtualization. The reason for simplifying the estimation of the overhead introduced by the virtualization technologies in eCAL through a factor $\gamma_{\mathrm{v}}$ is that the virtualization technologies in cloud systems are relatively well established while in communication systems they are still emerging. Furthermore, they depend on the configuration of the cloud-edge virtualization (e.g. bare-metal vs containers) and platform stacks (e.g. K8s vs K3s vs MicroK8s) as well as MLOps technology stack selection. For instance, there is significant difference in control-plane overhead on the platform layer of the CCRA depending on the number of control plane and worker plane nodes across K8s versions as shown in [57]. Furthermore, there is significant overhead difference between the standard O-RAN MLOps stack vs. NAOMI, a recently introduced distributed alternative [58]. In the future, as technologies will mature and consolidate, it is possible to set this factor to 0, tune the overhead in Eqs. (2) and incorporate compute and control plane overhead in Eq. (30) to enable more fine grained analytical study of virtualization impact.
@@
Finally, the proposed eCAL metric computing the corresponding energy consumption per bit over the lifecycle of the currently deployed AI model can be expressed as: